Write the slope-intercept equation for the line that passes through (5, -12) and (8, 12). Please show all of your work.
Answer by stanbon(75887)

Write the slope-intercept equation for the line that passes through (5, -12) and (8, 12). Please show all of your work.
---
slope = (12--12)/(8-5) = 24/3 = 8
----
Form: y = mx+b
Solve for "b":
12 = 8*8 + b
b = -52
----
Equation:
y = 8x - 52
========================
Cheers,
Stan H.
===============